OpenStudy (anonymous):

Simplify. 6 ² - (1/2) ³ ------------ 7/8 + 5/12 Show all work. This is what I got is it right? 6 ² - (1/2) ³ ------------ 7/8 + 5/12 36 - 1/8 -------- 31/24 24 [ 36 - 1/8 ] --------------- 31 24 [ 287/8 ] ------------ 31 3 [ 287 ] --------- 31 861 ---- 31 6 ² - (1/2) ³ ------------ 7/8 + 5/12 36 - 1/8 -------- 31/24 24 [ 36 - 1/8 ] --------------- 31 24 [ 287/8 ] ------------ 31 3 [ 287 ] --------- 31 861 ---- 31

OpenStudy (anonymous):

Sounds right to me. BTW - Congratulations. You are the first student who actually showed work and asked for confirmation whether it is right.

OpenStudy (anonymous):

I do most of the time.. Tonight has been an exception..LOL.. But most of the time I just come on here to check my answers.

OpenStudy (anonymous):

You are a rare bird here and thank you for that approach. Like I said, congratulations!

OpenStudy (anonymous):

Yes you are right. It is 861/31.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

\[\frac{6{}^{\wedge}2-\left(\frac{1}{2}\right)^3}{\frac{7}{8}+\frac{5}{12}}=\frac{36-\frac{1}{8}}{\frac{31}{24}}=\frac{\frac{287}{8}}{\frac{31}{24}}=\frac{861}{31} \]